A falcon, when diving, can travel five times as fast as a pheasant's top speed.If the total speeds for these two birds is 222 miles per hour, find the fastest speed of the falcon and the fastest speed of the pheasant.
Answer by checkley77(12844)

x+5x=222
6x=222
x=222/6
x=37 mph for the pheasant.
5*37=185 mph for the falcon.
Proof:
37+185=222
222=222